Find the correct value of Y to make a valid probability distributions


1
Expert's answer
2021-03-31T15:35:05-0400

Given Probability distribution is-


X - 1 2 3 4 5

P(x) - 0.11 0.15 0.35 Y 0.24


As we know "\sum P(x_i)=1"


"\Rightarrow 0.11+0.15+0.35+Y+0.24=1\\\Rightarrow Y+0.85=1\\\Rightarrow Y=1-0.85=0.15"